# Self-Managing DRAM: A Low-Cost Framework for Enabling Autonomous and Efficient DRAM Maintenance Operations

### v8

@@ 5.5 Other Maintenance Operation Improvements
SMD can seamlessly enable optimizations for the three maintenance operations we extensively describe (SMD-FR section 5.1, SMD-DRP section 5.2, and SMD-MS 5.3). At a high-level, any optimization idea or technique that is independent from the organization of multiple DRAM chips into a module or a rank (i.e., any idea that can be leveraged by some technique in a single DRAM chip) can be implemented in an SMD chip. Implementing such an idea or technique in an SMD chip does not require further changes to the DRAM interface.
@@ 6.2 DRAM Chip Modifications
DRAM Chip Modifications for Maintenance Mechanisms. Besides these changes that are the core of the SMD substrate, a particular maintenance mechanism may incur additional area overhead. We evaluate the DRAM chip area overhead of the three maintenance mechanisms presented in section 5.1, section 5.2, and section 5.3. The simple fixed-rate refresh mechanism, SMD-FR, requires only additional area in a DRAM chip. The DRAM chip area overhead of the refresh mechanism is less than 0.1% of a typical DRAM chip. SMD-DRP requires a large Counter Table with 1224 counters per bank for the RowHammer threshold value that we use in our performance evaluation. SMD-DRP requires area, which is 7.0% of a typical DRAM chip.171717Note that the area overhead comes especially from Graphene’s expensive CAM structures [61]. Cheaper RowHammer mitigation mechanisms can be implemented with SMD, as we describe in section 5.5. The control logic of SMD-MS is similar to the control logic of SMD-FR and it requires only additional area excluding the area of the ECC engine, which is already implemented by DRAM chips that support in-DRAM ECC.
@@ 6.3 Memory Controller Modifications
We slightly modify the MC’s scheduling mechanism to retry a rejected ACT command as we explain in section 4.3. Upon receiving an ACT_NACK, the MC marks the bank as precharged. An existing MC already implements control circuitry to pick an appropriate request from the request queue and issue the necessary DRAM command based on the DRAM bank state (e.g., ACT to a precharged bank or RD if the corresponding row is already open) by respecting the DRAM timing parameters. The ACT Retry Interval (ARI) is simply a new timing parameter that specifies the minimum time interval for issuing an ACT to a lock region after receiving an ACT_NACK from the same region. Therefore, SMD can be implemented in existing MCs with only slight modifications by leveraging the existing request scheduler (e.g., FR-FCFS [160]).
@@
Address Mapping Schemes. SMD does not require modifications to how a physical address is mapped to a memory-controller-visible DRAM address. In section 8, we evaluate SMD using a mapping scheme that interleaves consecutive cache blocks across channels, columns, ranks, banks, and rows, in that order. We also evaluate SMD using the mapping scheme described in [206] that aims to exploit bank-level parallelism. For this mapping scheme, SMD-FR provides 8.7% speedup over the baseline system, on average across 4c-high workloads. We leave detailed evaluation of SMD with different address mapping schemes for future work.
@@ 7 Experimental Methodology
We extend Ramulator [151, 150] to implement and evaluate the three SMD maintenance mechanisms (SMD-FR, SMD-DRP, and SMD-MS) that we describe in section 5.1, section 5.2, and section 5.3. We use DRAMPower [207, 208] to evaluate DRAM energy consumption. We use Ramulator in CPU-trace driven mode executing traces of representative sections of our workloads collected with a custom Pintool [209]. We warm-up the caches by fast-forwarding 100 million (M) instructions. We simulate each representative trace for 500M instructions (for multi-core simulations, until each core executes at least 500M instructions).
We use the system configuration provided in Table 1 in our evaluations. Although our evaluation is based on DDR4 DRAM, the modifications required to enable SMD can be adopted in other DRAM standards, as we explain in section 6.
@@
Comparison Points. We compare SMD to memory controller/DRAM co-design techniques. First, DARP [19] intelligently schedules per-bank refresh commands to idle banks while the memory controller is in write mode. Doing so reduces delays imposed by refresh operations on memory demand requests (e.g., load instructions executed by the processor). Second, DSARP [19] implements DARP and also modifies DRAM chip design and the DRAM interface to concurrently perform a refresh operation in one subarray and a memory access in another subarray, thereby hiding the latency of refresh operations and reducing delays they impose on memory requests. We modify DARP and DSARP such that they also reduce delays imposed on memory requests by Deterministic RowHammer Protection’s (section 5.2) victim row refresh operations.
Parameter Values of Maintenance Mechanisms and SMD. SMD-FR (section 5.1) refreshes a DRAM row every . SMD-FR refreshes 8 DRAM rows when it locks a DRAM region (RG = 8) with each refresh operation and postpones up to 8 such refresh operations (N = 8). Based on [20], for SMD-VR, we conservatively assume 0.1% of rows in each bank need to be refreshed every while the rest retain their data correctly for and more. SMD-VR uses a 8K-bit Bloom Filter with 6 hash functions. SMD-PRP refreshes the victims of an activated row with a high probability, i.e., = 1%. SMD-DRP (section 5.2) refreshes the victims before the aggressor is activated (512) times during a tREFW (). SMD-MS (section 5.3) operates with an aggressive 5-minute scrubbing period.191919We analyze SMD-DRP, and SMD-MS at various configurations. Their performance overheads are relatively low even at more aggressive settings. We configure SMD with 16 lock regions (section 4.2) in a DRAM bank unless stated otherwise. SMD uses an ACT Retry Interval () of (section 4.3).
@@ 8.1 Single-core Performance
First, SMD-FR provides 4.8% average speedup across all workloads. Although SMD-FR and DDR4 have similar average latency to refresh a single DRAM row, SMD-FR outperforms baseline DDR4 refresh due to two main reasons: SMD-FR 1) enables concurrent access and refresh of two different lock regions within a DRAM bank and 2) completely eliminates the REF commands on the DRAM command bus, thereby reducing the command bus utilization. Second, SMD-VR provides a higher 5.5% average speedup across all workloads as it mitigates the overhead of periodic refresh operations by refreshing different rows at different refresh rates. Third, SMD-FR + SMD-DRP (SMD-FR + SMD-PRP) provides 4.8% (4.2%) average speedup across all workloads. We observe that for a RowHammer threshold of 512 row activations (section 5.2), SMD-DRP and SMD-PRP can almost completely overlap the latency of victim row refresh operations in one lock region with the latency of a memory access to another lock region. Fourth, SMD-FR + SMD-MS provides 5.0% average speedup across all workloads. SMD enables low-overhead in-DRAM memory scrubbing for a relatively high memory scrubbing rate (each DRAM cell gets scrubbed every 5 minutes). The scrubbing operations induce negligible performance overhead: The average speedup provided by SMD-FR + SMD-MS is within <0.1% of SMD-FR if we exclude cond-mat-2003.212121SMD-FR + SMD-MS provides a relatively high 23.9% speedup for cond-mat-2003. SMD-FR provides a smaller speedup than SMD-Combined only for cond-mat-2003. For the same workload, SMD-FR has 5.1% higher average memory access latency than SMD-Combined. Even though SMD-Combined rejects (ACT_NACKs) more activate commands (15622) than SMD-FR (14286) on average across all four DRAM channels, in SMD-Combined, load instructions at the head of the processor core’s reorder buffer commit earlier on average than in SMD-FR. Consequently, SMD-Combined is less frequently bottlenecked by reorder buffer stalls (15.5% of execution time) than SMD-FR (40.2% of execution time). We attribute SMD-FR’s relatively poor performance for cond-mat-2003 to reorder buffer stalls. We hypothesize that the duration and the frequency of ACT_NACKs in SMD-FR affect main memory request scheduling in a way that older requests in the reorder buffer experience higher memory latency. Fifth, SMD-Combined provides 5.0% average speedup across all workloads. SMD-Combined provides 84.7% of the speedup provided by No-Refresh on average across all evaluated single-core workloads, while also improving system robustness with RowHammer protection and memory scrubbing (which No-Refresh does not do).
@@ 8.2 Multi-core Performance
First, SMD provides substantial performance improvements for memory intensive workloads with relatively high LLC MPKI (i.e., 4c-medium and 4c-high workloads). For example, SMD-Combined provides 5.1% (8.2%) and 8.9% (10.5%) average (maximum) speedup across all evaluated 4c-medium and 4c-high workloads, respectively. SMD does so by overlapping the latency of a majority of the maintenance operations with the latency of memory demand requests (e.g., load instructions). For 4c-high workloads, SMD-Combined provides 88.3% of the average speedup provided by No-Refresh even though SMD-Combined performs periodic refresh (section 5.1), RowHammer protection (section 5.2), and frequent memory scrubbing (section 5.3) that No-Refresh does not perform. Second, all evaluated SMD configurations provide similar speedups. For example, SMD-FR, SMD-FR + SMD-DRP, SMD-FR + SMD-MS, and SMD-Combined all provide approximately 8.9%; SMD-VR provides 9.2% and SMD-FR + SMD-PRP provides 7.4% average speedup across the evaluated 4c-high workloads. This indicates that SMD is able to effectively hide the latencies of RowHammer protection and memory scrubbing operations. Third, SMD provides small performance improvements for non-memory-intensive workloads with relatively low LLC MPKI (i.e., 4c-low workloads). For such workloads, completely eliminating maintenance operations (i.e., No-Refresh) also provides small performance improvements. SMD-Combined provides 88.8% of the speedup provided by No-Refresh for 4c-low workloads.
@@ 9.1 Prioritizing Memory Requests
“Pause Maintenance” Policy (SMD-PMP). The key idea of SMD-PMP is to pause an ongoing maintenance operation for a lock region when the memory controller issues an activate command to the locked region. The SMD chip could resume the maintenance operation when the memory controller precharges the bank (i.e., finishes accessing the lock region). For example, the SMD chip sequentially refreshes eight rows in a lock region (section 5.1) during a periodic refresh. If the chip receives an activate command to this lock region while only four out of eight rows have been refreshed, the chip does not continue refreshing the fifth row, but yields control of the lock region to the memory controller.
Latency of Pausing Maintenance Operations. Even if the SMD chip decides to pause the ongoing maintenance operation, the activate command cannot immediately proceed because the lock region might not be in the precharged state (i.e., the DRAM bank might not be ready to activate a row). The DRAM chip requires up to (section 2.1) to bring the under-maintenance lock region’s state to the precharged state (in case an activate command was just issued by the maintenance mechanism), depending on the lock region’s state when the activate command is received.
Estimating the Performance of SMD-PMP. We model SMD-PMP-FR (building on SMD-FR section 5.1) that refreshes only one DRAM row in a lock region before releasing the lock (i.e., SMD-PMP-FR locks a region for at most ). Fig. 12 shows normalized weighted speedup of 60 four-core workloads (20 per memory intensity level) with SMD-FR and SMD-PMP-FR on the left and the number of ACT_NACKs over the number of issued activate commands (which we call “ACT_NACK rate”), averaged across all workloads in every intensity level on the right. SMD-FR is configured as described in section 7 and it refreshes 8 DRAM rows in a lock region before releasing the lock.
@@ 9.2 Sensitivity to Number of Lock Regions
We conclude that SMD with 16 lock regions in a bank yields robust performance and DRAM energy improvements at relatively low DRAM chip area cost (see section 6).
@@ 10 Related Work
Changing the DRAM Interface. Several prior works [218, 219, 220, 221, 222] propose using high-speed serial links and packet-based protocols in DRAM chips. SMD differs from prior works in two key aspects. First, none of these works describe how to implement maintenance mechanisms completely within DRAM. We propose eight SMD-based maintenance mechanisms (section 5) and rigorously evaluate three (section 5.1, section 5.2, and section 5.3) of them. Second, prior works significantly modify or even overhaul the existing DRAM interface, which makes their proposals more difficult to adopt compared to SMD, which adds only a single ACT_NACK signal to the existing DRAM interface and requires slight modifications in the MC.
Mitigating DRAM Refresh Overheads. Many previous works [19, 20, 17, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46] propose techniques to reduce performance and energy overheads of DRAM refresh. SMD can seamlessly integrate many of these techniques along with other efficient maintenance operations (section 5).

### v9

@@ 5.5 Other Maintenance Operation Improvements
SMD can seamlessly enable optimizations for the three maintenance operations we extensively describe (SMD-FR section˜5.1, SMD-DRP section˜5.2, and SMD-MS 5.3). At a high-level, any optimization idea or technique that is independent from the organization of multiple DRAM chips into a module or a rank (i.e., any idea that can be leveraged by some technique in a single DRAM chip) can be implemented in an SMD chip. Implementing such an idea or technique in an SMD chip does not require further changes to the DRAM interface.
@@ 6.2 DRAM Chip Modifications
DRAM Chip Modifications for Maintenance Mechanisms. Besides these changes that are the core of the SMD substrate, a particular maintenance mechanism may incur additional area overhead. We evaluate the DRAM chip area overhead of the three maintenance mechanisms presented in section˜5.1, section˜5.2, and section˜5.3. The simple fixed-rate refresh mechanism, SMD-FR, requires only additional area in a DRAM chip. The DRAM chip area overhead of the refresh mechanism is less than 0.1% of a typical DRAM chip. SMD-DRP requires a large Counter Table with 1224 counters per bank for the RowHammer threshold value that we use in our performance evaluation. SMD-DRP requires area, which is 7.0% of a typical DRAM chip.171717Note that the area overhead comes especially from Graphene’s expensive CAM structures [61]. Cheaper RowHammer mitigation mechanisms can be implemented with SMD, as we describe in section 5.5. The control logic of SMD-MS is similar to the control logic of SMD-FR and it requires only additional area excluding the area of the ECC engine, which is already implemented by DRAM chips that support in-DRAM ECC.
@@ 6.3 Memory Controller Modifications
We slightly modify the MC’s scheduling mechanism to retry a rejected ACT command as we explain in section˜4.3. Upon receiving an ACT_NACK, the MC marks the bank as precharged. An existing MC already implements control circuitry to pick an appropriate request from the request queue and issue the necessary DRAM command based on the DRAM bank state (e.g., ACT to a precharged bank or RD if the corresponding row is already open) by respecting the DRAM timing parameters. The ACT Retry Interval (ARI) is simply a new timing parameter that specifies the minimum time interval for issuing an ACT to a lock region after receiving an ACT_NACK from the same region. Therefore, SMD can be implemented in existing MCs with only slight modifications by leveraging the existing request scheduler (e.g., FR-FCFS [160]).
@@
Address Mapping Schemes. SMD does not require modifications to how a physical address is mapped to a memory-controller-visible DRAM address. In section˜8, we evaluate SMD using a mapping scheme that interleaves consecutive cache blocks across channels, columns, ranks, banks, and rows, in that order. We also evaluate SMD using the mapping scheme described in [206] that aims to exploit bank-level parallelism. For this mapping scheme, SMD-FR provides 8.7% speedup over the baseline system, on average across 4c-high workloads. We leave detailed evaluation of SMD with different address mapping schemes for future work.
@@ 7 Experimental Methodology
We extend Ramulator [151, 150] to implement and evaluate the three SMD maintenance mechanisms (SMD-FR, SMD-DRP, and SMD-MS) that we describe in section˜5.1, section˜5.2, and section˜5.3. We use DRAMPower [207, 208] to evaluate DRAM energy consumption. We use Ramulator in CPU-trace driven mode executing traces of representative sections of our workloads collected with a custom Pintool [209]. We warm-up the caches by fast-forwarding 100 million (M) instructions. We simulate each representative trace for 500M instructions (for multi-core simulations, until each core executes at least 500M instructions).
We use the system configuration provided in Table 1 in our evaluations. Although our evaluation is based on DDR4 DRAM, the modifications required to enable SMD can be adopted in other DRAM standards, as we explain in section˜6.
@@
Comparison Points. We compare SMD to memory controller/DRAM co-design techniques. First, DARP [19] intelligently schedules per-bank refresh commands to idle banks while the memory controller is in write mode. Doing so reduces delays imposed by refresh operations on memory demand requests (e.g., load instructions executed by the processor). Second, DSARP [19] implements DARP and also modifies DRAM chip design and the DRAM interface to concurrently perform a refresh operation in one subarray and a memory access in another subarray, thereby hiding the latency of refresh operations and reducing delays they impose on memory requests. We modify DARP and DSARP such that they also reduce delays imposed on memory requests by Deterministic RowHammer Protection’s (section˜5.2) victim row refresh operations.
Parameter Values of Maintenance Mechanisms and SMD. SMD-FR (section˜5.1) refreshes a DRAM row every . SMD-FR refreshes 8 DRAM rows when it locks a DRAM region (RG = 8) with each refresh operation and postpones up to 8 such refresh operations (N = 8). Based on [20], for SMD-VR, we conservatively assume 0.1% of rows in each bank need to be refreshed every while the rest retain their data correctly for and more. SMD-VR uses a 8K-bit Bloom Filter with 6 hash functions. SMD-PRP refreshes the victims of an activated row with a high probability, i.e., = 1%. SMD-DRP (section˜5.2) refreshes the victims before the aggressor is activated (512) times during a tREFW (). SMD-MS (section˜5.3) operates with an aggressive 5-minute scrubbing period.191919We analyze SMD-DRP, and SMD-MS at various configurations. Their performance overheads are relatively low even at more aggressive settings. We configure SMD with 16 lock regions (section˜4.2) in a DRAM bank unless stated otherwise. SMD uses an ACT Retry Interval () of (section˜4.3).
@@ 8.1 Single-core Performance
First, SMD-FR provides 4.8% average speedup across all workloads. Although SMD-FR and DDR4 have similar average latency to refresh a single DRAM row, SMD-FR outperforms baseline DDR4 refresh due to two main reasons: SMD-FR 1) enables concurrent access and refresh of two different lock regions within a DRAM bank and 2) completely eliminates the REF commands on the DRAM command bus, thereby reducing the command bus utilization. Second, SMD-VR provides a higher 5.5% average speedup across all workloads as it mitigates the overhead of periodic refresh operations by refreshing different rows at different refresh rates. Third, SMD-FR + SMD-DRP (SMD-FR + SMD-PRP) provides 4.8% (4.2%) average speedup across all workloads. We observe that for a RowHammer threshold of 512 row activations (section˜5.2), SMD-DRP and SMD-PRP can almost completely overlap the latency of victim row refresh operations in one lock region with the latency of a memory access to another lock region. Fourth, SMD-FR + SMD-MS provides 5.0% average speedup across all workloads. SMD enables low-overhead in-DRAM memory scrubbing for a relatively high memory scrubbing rate (each DRAM cell gets scrubbed every 5 minutes). The scrubbing operations induce negligible performance overhead: The average speedup provided by SMD-FR + SMD-MS is within <0.1% of SMD-FR if we exclude cond-mat-2003.212121SMD-FR + SMD-MS provides a relatively high 23.9% speedup for cond-mat-2003. SMD-FR provides a smaller speedup than SMD-Combined only for cond-mat-2003. For the same workload, SMD-FR has 5.1% higher average memory access latency than SMD-Combined. Even though SMD-Combined rejects (ACT_NACKs) more activate commands (15622) than SMD-FR (14286) on average across all four DRAM channels, in SMD-Combined, load instructions at the head of the processor core’s reorder buffer commit earlier on average than in SMD-FR. Consequently, SMD-Combined is less frequently bottlenecked by reorder buffer stalls (15.5% of execution time) than SMD-FR (40.2% of execution time). We attribute SMD-FR’s relatively poor performance for cond-mat-2003 to reorder buffer stalls. We hypothesize that the duration and the frequency of ACT_NACKs in SMD-FR affect main memory request scheduling in a way that older requests in the reorder buffer experience higher memory latency. Fifth, SMD-Combined provides 5.0% average speedup across all workloads. SMD-Combined provides 84.7% of the speedup provided by No-Refresh on average across all evaluated single-core workloads, while also improving system robustness with RowHammer protection and memory scrubbing (which No-Refresh does not do).
@@ 8.2 Multi-core Performance
First, SMD provides substantial performance improvements for memory intensive workloads with relatively high LLC MPKI (i.e., 4c-medium and 4c-high workloads). For example, SMD-Combined provides 5.1% (8.2%) and 8.9% (10.5%) average (maximum) speedup across all evaluated 4c-medium and 4c-high workloads, respectively. SMD does so by overlapping the latency of a majority of the maintenance operations with the latency of memory demand requests (e.g., load instructions). For 4c-high workloads, SMD-Combined provides 88.3% of the average speedup provided by No-Refresh even though SMD-Combined performs periodic refresh (section˜5.1), RowHammer protection (section˜5.2), and frequent memory scrubbing (section˜5.3) that No-Refresh does not perform. Second, all evaluated SMD configurations provide similar speedups. For example, SMD-FR, SMD-FR + SMD-DRP, SMD-FR + SMD-MS, and SMD-Combined all provide approximately 8.9%; SMD-VR provides 9.2% and SMD-FR + SMD-PRP provides 7.4% average speedup across the evaluated 4c-high workloads. This indicates that SMD is able to effectively hide the latencies of RowHammer protection and memory scrubbing operations. Third, SMD provides small performance improvements for non-memory-intensive workloads with relatively low LLC MPKI (i.e., 4c-low workloads). For such workloads, completely eliminating maintenance operations (i.e., No-Refresh) also provides small performance improvements. SMD-Combined provides 88.8% of the speedup provided by No-Refresh for 4c-low workloads.
@@ 9.1 Prioritizing Memory Requests
“Pause Maintenance” Policy (SMD-PMP). The key idea of SMD-PMP is to pause an ongoing maintenance operation for a lock region when the memory controller issues an activate command to the locked region. The SMD chip could resume the maintenance operation when the memory controller precharges the bank (i.e., finishes accessing the lock region). For example, the SMD chip sequentially refreshes eight rows in a lock region (section˜5.1) during a periodic refresh. If the chip receives an activate command to this lock region while only four out of eight rows have been refreshed, the chip does not continue refreshing the fifth row, but yields control of the lock region to the memory controller.
Latency of Pausing Maintenance Operations. Even if the SMD chip decides to pause the ongoing maintenance operation, the activate command cannot immediately proceed because the lock region might not be in the precharged state (i.e., the DRAM bank might not be ready to activate a row). The DRAM chip requires up to (section˜2.1) to bring the under-maintenance lock region’s state to the precharged state (in case an activate command was just issued by the maintenance mechanism), depending on the lock region’s state when the activate command is received.
Estimating the Performance of SMD-PMP. We model SMD-PMP-FR (building on SMD-FR section˜5.1) that refreshes only one DRAM row in a lock region before releasing the lock (i.e., SMD-PMP-FR locks a region for at most ). Fig. 12 shows normalized weighted speedup of 60 four-core workloads (20 per memory intensity level) with SMD-FR and SMD-PMP-FR on the left and the number of ACT_NACKs over the number of issued activate commands (which we call “ACT_NACK rate”), averaged across all workloads in every intensity level on the right. SMD-FR is configured as described in section˜7 and it refreshes 8 DRAM rows in a lock region before releasing the lock.
@@ 9.2 Sensitivity to Number of Lock Regions
We conclude that SMD with 16 lock regions in a bank yields robust performance and DRAM energy improvements at relatively low DRAM chip area cost (see section˜6).
@@ 9.8 SMD-based Probabilistic
9.9 ACT_NACK Divergence Across Chips
In section˜4.4, we explain divergence in SMD maintenance operations can happen when different DRAM chips in the same rank perform maintenance operation at different times. Such a divergence leads to a partial row activation when the activated row is in a locked region in some DRAM chips but not in others. To handle partial row activations, we develop three policies.
In Fig. 18, we compare the performance and energy savings of SMD-FR and SMD-VR when using the three ACT_NACK divergence handling policies across 4c-high workloads for tREFW = . The plots show results for the common-case (CC) and worst-case (WC) scenarios with regard to when maintenance operations happen across different SMD chips in the same rank. In the common-case scenario, the DRAM chips generally refresh the same row at the same time due to sharing the same DRAM architecture design. However, refresh operations in some of the DRAM chips may still diverge during operation depending on the refresh mechanism that is in use. For example, SMD-VR refreshes retention-weak rows, whose locations may differ across the DRAM chips, at a higher rate compared to other rows, resulting in divergence in refresh operations across the DRAM chips in a rank. In the worst-case scenario, we deliberately configure the DRAM chips to refresh different rows at different times. For this, we 1) delay the first refresh operation in by , where and is the latency of a single refresh operation, and 2) set the Lock Region Counter (LRC) of to .
We make three key observations. First, the performance and DRAM energy consumption only slightly varies across different divergence handling policies in both common-case and worst-case scenarios. This is because, after a partial row activation happens, a different row that is not in a locked region in all off of the DRAM chips often does not exist in the memory request queue. As a result, the Precharge and Hybrid policies perform similarly to the Wait policy.
Second, SMD-FR performs worse than the DDR4 baseline in the worst-case refresh distribution scenario (i.e., average slowdown of 2.6% with the Wait policy). Certain individual workloads experience even higher slowdown (up to 12.8% with the Wait policy). The reason is that, when different DRAM chips lock a region at different times, the total duration during which the lock region is unavailable becomes times the duration when all chips simultaneously refresh the same lock region. This significantly increases the performance overhead of refresh operations.
Third, SMD-VR does not suffer much from the divergence problem and outperforms the DDR4 baseline even in the worst-case scenario. This is because SMD-VR mitigates the DRAM refresh overhead by significantly reducing the number of total refresh operations by exploiting retention-time variation across DRAM rows. Thus, the benefits of eliminating many unnecessary refresh operations surpass the overhead of ACT_NACK divergence.
We conclude that 1) although SMD-FR suffers from noticeable slowdown for the worst-case scenario, which should not occur in a well-designed system, it still provides comparable performance to conventional DDR4 and 2) SMD-VR outperforms the baseline and saves DRAM energy even in the worst-case scenario.
@@ 10 Related Work
Changing the DRAM Interface. Several prior works [218, 219, 220, 221, 222] propose using high-speed serial links and packet-based protocols in DRAM chips. SMD differs from prior works in two key aspects. First, none of these works describe how to implement maintenance mechanisms completely within DRAM. We propose eight SMD-based maintenance mechanisms (section˜5) and rigorously evaluate three (section˜5.1, section˜5.2, and section˜5.3) of them. Second, prior works significantly modify or even overhaul the existing DRAM interface, which makes their proposals more difficult to adopt compared to SMD, which adds only a single ACT_NACK signal to the existing DRAM interface and requires slight modifications in the MC.
Mitigating DRAM Refresh Overheads. Many previous works [19, 20, 17, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46] propose techniques to reduce performance and energy overheads of DRAM refresh. SMD can seamlessly integrate many of these techniques along with other efficient maintenance operations (section˜5).